# ZeroCAP: Zero-Shot Multi-Robot Context Aware Pattern Formation via Large Language Models

### v1

@@ IV-C LLM-based Pattern Former
Upon accurately delineating the geometry of $O$, the context-aware multi-robot deployment system advances to determining goal position coordinates using a LLM. This stage involves the LLM being provided with the number of available robots $R$, interpreting the Pattern Formation Instruction $PI$ alongside the object’s spatial characteristics, represented through edges $E$ and vertices $V$ to provide an output representing positions for robots.
The LLM’s task is to integrate the $PI$ with the spatial data of $E$ and $V$ to determine the optimal pattern formation. This involves comprehensive reasoning to align the robots according to $PI$, while considering the shape of the object and the available robotic agents $R$.
In particular, our methodology eschews the direct use of the entire Masked Image $I_{m}$ as input to the LLM, a decision underscored by considerations of computational efficiency and token costs. Direct image is resource-intensive, whereas leveraging $E$ and $V$ markedly reduces computational load especially when accounting for tokens in the context of LLM’s. This choice highlights our system’s capacity for real-time processing and deployment efficiency while also remaining inexpensive in computation and practicality.
The culmination of this process is the production of a coordinate list $C_{xy}$ as seen in Line 6 of Algorithm 1, which prescribes exact deployment coordinates for each robot.
@@ IV-D Robot Deployment
Upon generating the deployment coordinates $C_{xy}$ with the LLM, the ZeroCAP system initiates robot deployment, translating algorithmic output into coordinated physical actions.
Each robot $r_{i},i\in N$ is directed toward its assigned location $(x_{i},y_{i})$ under centralized control, ensuring precise navigation. Upon arrival, the system verifies each robot’s position against $C_{xy}$, adjusting as necessary to rectify any positional inaccuracies, thereby ensuring the formation aligns perfectly with the pattern instructions. This is delineated in Lines 7 to 9 and 14 to 16 in Algorithm 1.
The successful deployment of robots in the context of the Object of Interest $O$ marks the completion of the pattern formation task, embodying the ZeroCAP system’s capability to execute context-driven, instruction-based patterns through autonomous robotic agents.
@@ V Experiment
(a) General pattern formation tasks: robots arrange themselves around the object of interest based on user specification
(b) Infill pattern formation tasks: robots occupy the interior space of a region based on user specification
(c) Caging formation tasks: creating a cage like boundary around the object
(d) Application oriented task: surrounding the incorrectly parked car with four robots.
Figure 4: Illustrations of the three types of pattern formation tasks and a task demonstration of Fig 1, showcasing the system’s adaptability to various instruction types and object configurations. In all tasks, robots are represented as black points initially distributed at the bottom of the scene in the initial environment, transitioning to form user-directed patterns in the final environment. This illustrates the dynamic capability of our system to interpret instructions and spatially reorganize the robots accordingly. Please zoom in on the images for details.
Our experimental evaluation investigates the ZeroCAP system’s zero-shot pattern formation capabilities across ten tasks as shown in Fig. 4, delineated into three categories as prefaced in Section I: general pattern formations as shown in subfigure 4(a), infill pattern formations as shown in subfigure 4(b), and caging formation tasks as shown in subfigure 4(c) and the task referred to in Fig 1 as shown in subfigure 4(d).
Tasks are diversified by object setup to evaluate the system’s adaptability and discernment capabilities under varying environmental complexities. In the Single Object setup (ENV a, b, d, g), the environment contains a solitary object, simplifying the identification process. The Multi-Object setup (ENV c, e, h) introduces additional objects as distractors, challenging the system to isolate the target object based on explicit user instructions. Lastly, the Hidden Object setup (ENV f, i, park) presents a more complex scenario where the natural language instruction implies the object of interest without explicit mention, compelling the system to employ reasoning to deduce the correct object.
@@ V-A Baseline Comparisons
Our system’s performance was benchmarked against three baselines. The first is an RL-based simplified approach with hand-crafted rewards, serving primarily as a pattern-formation benchmark without natural language comprehension. The second and third baselines use GPT4 Vision [33] and llava-v1.6 [34], respectively, relying solely on vision-language models to complete the tasks. The LLMs employed in our experiments are GPT-4 [35], llama70b [36], and Claude [37], contrasting their performance against each other and the baselines. For ablation studies, the default configuration of the ZeroCAP system utilizes GPT-4 as the primary LLM.
@@ V-B Environment
Experiments were conducted in a 2D simulation using Pygame, with robots as point-mass entities capable of precise movements to target coordinates. The setup includes multi-object tasks with infill and caging requirements, omitting obstacle avoidance behaviors to focus on the pattern formation tasks. A subset of tasks was also executed on a real testbed, featuring Hamster mobile robots [38] under an overhead camera setup. Fig. 5 showcases two tasks executed in a real-world setting, serving as a tangible proof of concept that corroborates our simulation results. To facilitate these tasks, the environment is simplified into a grid-based image, with each grid box approximately matching the size of a robot. This adaptation allows the robots to retain their point mass entity characterization within the practical constraints of the real-world experiments, ensuring consistency in their deployment and movement. For demonstrations of simulator and real-world tasks, please refer to the supplementary video.
@@ V-C Evaluation Metrics
We employed two main evaluation metrics: Success Rates (SR) and Goal Condition Recall (GCR). SR indicates whether the robots formed the pattern successfully, verified by handcrafted ground truth labels, while GCR quantifies the proportion of robots correctly positioned within the pattern. A GCR of 1 implies a perfect match with the ground truth and, consequently, an SR of 1. Each task underwent 10 trials, with averages reported for SR and GCR, chosen to account for the inherent variability in LLM performance.
@@ VI Results and Analysis
TABLE I: Comparative analysis of pattern formation methodologies across General, Infill, and Caging tasks, evaluating Success Rates (SR) and Goal Condition Recall (GCR) metrics across different methods.
@@
ZeroCAP (lama-2-70b) 0.2 0.47 0.39 0.71 0.32 0.67 0.3 0.61
ZeroCAP (claude-3-opus) 0.57 0.78 0.72 0.88 0.64 0.73 0.64 0.79
Table I shows a comprehensive evaluation of different methodologies tested in this paper, revealing the effectiveness of the ZeroCAP system in comparison to baselines. The initial baselines, GPT-4 Vision and LLava v1.6, were employed to evaluate the feasibility of using VLMs directly for zero-shot context-aware pattern formation. The performance of these baselines was significantly poor. GPT-4 Vision generally failed to generate any coordinates, reflecting its known limitations with spatial localization tasks. LLava, while capable of producing coordinates, generated positions that were largely inaccurate and contextually irrelevant. These outcomes effectively demonstrate that direct application of VLMs for context-aware pattern formation is non-trivial, and without additional processing or reasoning mechanisms, VLMs fall short of providing viable solutions. The RL baseline exhibited a mixed performance across the task spectrum. It notably struggled with the general pattern formation tasks, where user specifications varied greatly and required a high degree of interpretation and flexibility. The RL method’s rigidity and lack of language comprehension rendered it unable to adapt to the nuanced demands of these tasks. On the other hand, the RL approach showed considerable success in Caging tasks, aligning with the more straightforward objective of these tasks where the overarching goal is consistently to form a barrier around the object.
Figure 5: Two real-world scenarios demonstrating ZeroCAP system’s ability to execute context-aware pattern formations.
The ZeroCAP methods, incorporating LLMs like GPT-4, llama-2-70b, and claude-3-opus, present a contrasting picture. The GPT-4 based ZeroCAP achieves the highest overall SR and GCR across all types of tasks, indicating its superior pattern recognition and execution in zero-shot conditions. This suggests that GPT-4’s advanced natural language processing abilities are effectively leveraged by the ZeroCAP system to interpret instructions and generate appropriate robotic behaviors. The performance of ZeroCAP using llama-2-70b and claude-3-opus varies. While the llama-2-70b variant has some success, its inconsistency across tasks points to possible limitations in its ability to handle complex spatial reasoning as compared to GPT-4. Claude-3-opus shows a relatively more consistent performance. Across different pattern formation types, ZeroCAP methods behave in a manner that suggests certain patterns may be more conducive to LLM-driven execution. The General pattern formation tasks appear to be the most challenging, possibly due to the abstract nature of the instructions and the need for more nuanced interpretation. Infill tasks, which involve a more defined area and clear objectives being in most cases to deploy robots inside a region, show the highest SR and GCR for GPT-4, indicating a strong alignment between the method’s capabilities and the task requirements. Caging tasks, while more structured, still present a challenge, but ZeroCAP methods demonstrate reasonable effectiveness.
@@ VI-A Ablation: Segmentation Methods
The ablation study presented in Table II examines the efficacy of CLIPSeg [5] and LangSAM as object segmentation methods within the context of language-prompted object identification. These methods, while capable of segmenting objects based on language instructions, do not possess the same reasoning abilities provided by full-fledged VLMs. The table compares the performance of these segmenters used in isolation (denoted by ‘Only’) to their performance when integrated into the ZeroCAP framework.
@@
It is evident that using CLIPSeg and LangSAM alone for object identification and segmentation results in modest success rates of 0.51 and 0.43, respectively. This suggests that while they can interpret natural language to some extent and perform segmentation, the lack of deeper contextual understanding and spatial reasoning affects their overall effectiveness in pattern formation tasks. This is especially a factor in case of hidden object task scenarios.
When these segmentation methods are incorporated into the ZeroCAP system, the success rates improve significantly to 0.78 for CLIPSeg and 0.77 for LangSAM. The integration of these segmenters within ZeroCAP likely benefits from additional contextual processing and spatial reasoning capabilities provided by the LLMs in the system, leading to more successful task outcomes.
@@ VI-B Ablation: Shape Descriptors
The ablation study presented in Table III evaluates the influence of different shape descriptors on the performance of the ZeroCAP system. ‘Increase in Tokens’ indicates the relative percentage increase in the number of tokens used by the LLM compared to using only ‘Edges’ as the shape descriptor, which serves as the baseline in this context.
@@
Analyzing the data, it is evident that using only ‘Edges’ as a shape descriptor achieves the highest SR of 0.77. This suggests that ‘Edges’ alone provide ample information for the LLM to accurately generate the desired patterns without requiring additional token input. In contrast, when ‘Edges and Vertices’ are combined, there is a modest increase in tokens by 2%, but the SR significantly decreases to 0.48. This decrease could imply that including vertices introduces a level of complexity that the LLM does not effectively translate into better pattern formation. Moreover, vertices as a descriptor could potentially be misleading for non-polygonal shapes where the definition of vertices is less clear, potentially leading to inaccurate pattern formation by the LLM.
The binary image matrix descriptor is obtained by converting the segmented object mask into a binary matrix, which is then resized to a 100x100 grid to fit within the token constraints of the LLM. This matrix serves as a input for the LLM, representing the shape in a form that can be directly processed. The LLM utilizes this grid-based representation to understand the spatial arrangement and generate target coordinates for the robots. The binary image matrix descriptor, incurs a 400% increase in token usage. This upsurge poses practicality concerns due to token limits and practical costs. Additionally, resizing may obscure vital shape details, affecting pattern accuracy. Despite a reasonable SR of 0.71, the trade-off between detail richness and token efficiency is crucial for the ZeroCAP system’s viability.
VII Discussion
The research presented in this paper marks a significant stride in the domain of multi-robot systems, particularly in the application of LLMs for zero-shot context aware pattern formation. Nonetheless, the current scope is confined to static patterns, leaving the exploration of dynamic patterns - those that evolve over time in response to environmental changes or ongoing tasks - as a promising avenue for future research. Dynamic pattern formation would not only extend the flexibility and adaptability of multi-robot systems but also increase their applicability in real-world scenarios such as search and rescue operations or environmental monitoring, where conditions are continually shifting.
Our current method’s reliance on centralized processing poses potential bottlenecks and failure points in scenarios requiring quick robot adaptability. Exploring decentralized or hybrid control systems could improve robustness and scalability by merging centralized planning with decentralized execution. Additionally, incorporating sensory feedback mechanisms may enhance decision-making, maintaining system effectiveness amid real-time environmental changes.
Our research lays a foundation for future advancements in robotic pattern formation and practical applications, extending from complex coordination tasks in surveillance, logistics, and warehousing to context-aware strategies. Progressing beyond current limitations and enhancing our foundational framework is vital for fully realizing robotics’ capabilities in intricate, context-sensitive environments.
VIII Conclusion
Our research has successfully introduced ZeroCAP, an innovative framework that harmonizes the advanced linguistic comprehension of large language models with the physical capabilities of robotic systems to achieve zero-shot pattern formation. The system’s adeptness at deciphering natural language instructions and materializing these into precise spatial arrangements has been demonstrated across various contexts, laying the groundwork for future advancements in multi-robot coordination. While our study focused on static patterns and centralized control, the insights gained provide a stepping stone for exploring dynamic pattern formations. ZeroCAP’s performance in context-aware task execution signal a promising direction for the integration of AI and robotics, propelling us towards more intelligent, adaptable, and efficient multi-robot applications in diverse real-world settings.

### v2

@@ IV-C LLM-based Pattern Former
After delineating the geometry of $O$, the system advances to generating goal position coordinates using a LLM represented as box ‘Sec IV. C’ in Fig. 2. The LLM receives the number of available robots $R$, $PI$, and the spatial characteristics of the object, represented by its edges $E$ and vertices $V$, to compute the optimal robot deployment positions. Instead of using the entire Masked Image $I_{m}$, we opt for the more computationally efficient use of $E$ and $V$, significantly reducing token overhead for LLMs and enhancing real-time performance. This approach ensures faster processing while keeping computational costs low.
The LLM integrates the $PI$ with the spatial data $E$ and $V$ to determine the optimal pattern formation. This involves comprehensive reasoning to align the robots according to $PI$, while considering the shape of the object and the available robotic agents $R$. The output of this stage is a list of coordinates $C_{xy}$ specifying precise deployment locations for each robot, as outlined in Line 6 of Algorithm 1.
@@ IV-D Robot Deployment
Once the deployment coordinates $C_{xy}$ are generated, the ZeroCAP system initiates robot deployment, translating these coordinates into coordinated actions.
Each robot $r_{i},i\in N$ is directed toward its assigned location $(x_{i},y_{i})$. Upon arrival, the system verifies each robot’s position against $C_{xy}$, making adjustments as needed to correct any deviations, ensuring the formation aligns with the pattern instructions. This is shown in Lines 7 to 8 and 13 to 15 in Algorithm 1.
The successful deployment of robots around the $O$ completes the pattern formation task, demonstrating ZeroCAP’s ability to execute context-aware, instruction-based patterns, without the need for any prior training.
@@ V Experiment
Figure 4: Illustrations of five real-world tasks performed by the ZeroCAP system, showcasing different pattern formations: (a) and (b) illustrate general pattern formations; (c) and (d) depict infill pattern formations; (e) demonstrates a caging task. Each task is executed based on natural language instructions provided at the top of each subfigure. Blue boundaries around the robots indicate their initial positions, while green boundaries represent their final positions. Please zoom in on the images for details.
We evaluate the ZeroCAP system’s zero-shot pattern formation capabilities across five real-world tasks, as shown in Fig. 4 and ten tasks performed in simulated environments. These tasks are categorized into three types: 1) General pattern formation, where robots arrange around the object of interest; 2) Infill pattern formation, where robots occupy the interior of a region; and 3) Caging formation, where robots create a boundary around the object. This categorization shifts from standard geometric patterns to contextually relevant formations, driven by natural language instructions and environmental cues, demonstrating the system’s flexibility.
The tasks also vary by object setup to test adaptability. In the Single Object setup as shown in Fig. 4, a single object simplifies identification. The Multi-Object setup, shown in Fig. 5 adds distractors, requiring the system to isolate the target from several options based on explicit instructions. In the Hidden Object setup, shown in Fig. 5 and Fig. 6, the target is implied rather than stated, requiring the system to reason based on contextual clues.
@@ V-A Baseline Comparisons
The system is benchmarked against three baselines. The first is an RL-based approach with hand-crafted rewards, focusing on pattern formation without natural language comprehension. The second and third baselines utilize GPT-4 Vision [34] and LLava-v1.6 [35], relying solely on vision-language models to complete tasks. The LLMs used in our experiments include GPT-4 [36], llama70b [37], and Claude [38], allowing for performance comparison across models and baselines. In ablation studies, GPT-4 is the default LLM configuration for ZeroCAP.
@@ V-B Environment
Experiments were conducted in both real-world settings and in 2D simulations. Real-world experiments used a testbed with Hamster mobile robots under an overhead camera [39], with the environment simplified into a grid-based image to ensure consistency in point-mass robot deployment. Fig. 4 highlights the real-world tasks, demonstrating the practical viability of our approach. Additionally, 2D simulations were carried out using the Pygame environment, as shown in Fig. 5, where robots were modeled as point-mass entities and moved precisely to target coordinates. The simulation setup included multi-object tasks with infill and caging requirements, excluding obstacle avoidance to focus solely on pattern formation. These simulation experiments served to rigorously validate the proposed methods across varied conditions and object setups. For more detailed demonstrations of the system under various setups and task categories, please refer to the supplementary video.
@@ V-C Evaluation Metrics
We used two evaluation metrics: Success Rate (SR) and Goal Condition Recall (GCR). SR indicates if the robots successfully formed the pattern by comparing final positions to handcrafted ground truth labels, with success requiring all positions to match. GCR measures the proportion of correctly positioned robots, with a GCR of 1 implying an SR of 1. Each task was run over 10 trials, and average SR and GCR were reported to account for LLM variability.
@@ VI Results and Analysis
Table I summarizes the evaluation of the tested methodologies, highlighting ZeroCAP’s effectiveness over the baselines. GPT-4 Vision and LLava v1.6, are tested for direct VLM-only zero-shot pattern formation, both performed poorly. GPT-4 Vision failed to generate meaningful coordinates, revealing its limitations in spatial reasoning. LLava produced inaccurate or irrelevant positions. These results show that using VLMs alone for context-aware pattern formation is inadequate without enhanced spatial reasoning capabilities. The RL-based baseline showed mixed results, struggling with general and infill pattern formation due to varied user instructions requiring flexibility. RL’s reliance on predefined rewards limits generalizability; while effective in task-specific training, it underperforms with generalized reward structures. We chose a generalized reward approach to enhance scalability, as designing specific rewards for every new task is impractical. Consequently, RL performed poorly in general and infill tasks but excelled in caging tasks, where the consistent goal of forming a barrier aligned with the reward design. This highlights RL’s strength in structured tasks but its limitations in adapting to nuanced, context-driven user instructions.
In contrast, the ZeroCAP system, leveraging LLMs like GPT-4, llama-2-70b, and Claude-3-opus, significantly outperformed the baselines. GPT-4-based ZeroCAP achieved the highest SR and GCR across all tasks, largely due to its superior language processing, allowing for better interpretation of complex instructions. The ZeroCAP system’s architecture, which decouples spatial reasoning from the VLM and utilizes edge representations, mitigated GPT-4 Vision’s limitations in direct spatial localization. Llama-2-70b performed moderately well but showed inconsistencies, particularly in tasks with hidden object scenarios requiring complex spatial reasoning. Claude-3-opus, while slightly less accurate than GPT-4, demonstrated more consistency across different task types.
Figure 5: Two simulation scenarios demonstrating ZeroCAP’s ability to execute context-aware pattern formations. Sim1 (left) depicts a caging task in a hidden object setup, where the object of interest is not directly specified and must be inferred. Sim2 (right) illustrates an infill task in a multi-object setup. Please zoom in on the images for details.
Figure 6: Illustration of a general pattern formation task as shown in concept Fig. 1, featuring a hidden object setup where the system must infer the wrongly parked car and deploy robots, represented by black points, to surround it.
Task-wise, general pattern formation proved the most challenging, likely due to the abstract nature of the instructions, requiring nuanced interpretation. Infill tasks, with clearer objectives, achieved the highest SR and GCR, particularly with GPT-4, as they mainly involved placing robots within a specified region. Caging tasks, although more structured, still posed difficulties, but ZeroCAP managed consistent barrier formations effectively.
TABLE I: Comparative analysis of pattern formation methods for General, Infill, and Caging tasks, evaluating Success Rates (SR) and Goal Condition Recall (GCR) metrics across different methods. Total represents the combined performance across all task categories. Bold numbers indicate the best performance among the tested methods.
@@
ZeroCAP (llama-2-70b) 0.39 0.47 0.55 0.71 0.62 0.74 0.52 0.64
ZeroCAP (claude-3-opus) 0.63 0.78 0.79 0.88 0.69 0.79 0.70 0.81
@@ VI-A Ablation: Segmentation Methods
Table II presents the results of an ablation study comparing the performance of CLIPSeg [5] and LangSAM as standalone segmentation methods and their integration within the ZeroCAP framework. While CLIPSeg and LangSAM achieve modest success rates (0.51 and 0.43, respectively) when used alone, their performance is limited by a lack of spatial reasoning and deeper contextual understanding, particularly in tasks requiring hidden object identification.
@@
When integrated into ZeroCAP, success rates rise to 0.78 and 0.77, respectively. This improvement demonstrates the added value of incorporating VLMs into the system, providing the necessary contextual reasoning and spatial capabilities to enhance overall performance.
@@ VI-B Ablation: Shape Descriptors
The ablation study presented in Table III examines the impact of different shape descriptors on ZeroCAP’s performance. ‘Increase in Tokens’ reflects the percentage increase in token usage compared to using only ‘Edges’ as the shape descriptor.
@@
Results show that using ‘Edges’ alone yields the highest SR (0.77), demonstrating that this descriptor provides sufficient information for the LLM without increasing token load. Adding ‘Vertices’ results in only a 2% token increase, but the SR drops to 0.48, indicating that the complexity introduced by vertices can hinder performance, especially for non-polygonal shapes where vertex definitions are less clear. The binary image matrix, created by converting and resizing the object mask to a $100\times 100$ grid, increases token usage by 400%, making it less practical. While it achieves a decent SR of 0.71, the token cost and potential loss of important shape features during resizing highlight the trade-off between detail and efficiency in ZeroCAP’s processing.
Overall, the results confirm that VLMs alone are insufficient for zero-shot context-aware pattern formation, underscoring the need for ZeroCAP. ZeroCAP outperforms baselines, particularly in tasks requiring nuanced reasoning. Ablation studies show that edges as shape descriptors yield the best success with minimal token load, highlighting ZeroCAP’s strength in integrating LLMs for robust, zero-shot pattern formation and setting a foundation for future multi-robot coordination research.
VII Conclusion
This work presents ZeroCAP, a framework integrating LLMs with MRS for zero-shot, context-aware pattern formation. By translating natural language instructions into precise spatial arrangements, ZeroCAP overcomes VLM spatial localization limitations by decoupling spatial reasoning from visual processing, using edge and vertex representations for more accurate pattern formation.
Currently, due to VLMs limitations in processing 3D data, being trained mostly on 2D images [11], this work is constrained to a 2D framework to ensure robustness. As VLMs evolve, future exploration could extend to 3D. The system is deployable with a map or overhead camera and could be expanded using multiple cameras in the future for wider scale and coverage. While our focus has been on static formations, dynamic formations that adapt to real-time changes hold significant promise. Further investigation of complex environments with partial occlusions could also improve robustness. This early-stage research lays the groundwork for more intuitive robotic coordination with implications in applications such as surveillance and logistics.
Acknowledgement
We would like to thank Dr. Tamzidul Mina at Sandia National Laboratories for his valuable feedback, which significantly enhanced the clarity and quality of this paper.